Q1 IF F(X)=ux+v .the number of points in ( u,v) ;f(x) =fof(X) for all x elongs to real number. is;   a) finitely many     b) infinitely many    c) 2   d) 4


Q2 LET F;(E,INFINITY) --R BE DEFINED BY F(x) =LOG(LOG(LOGX)) THEN THE CODOMAIN OF F IS ;

Grade:12

1 Answers

suchintan mondal
14 years ago

Q1. ans> fof(x)=uf(x)+v=u2x+uv+v

now f(x)=fof(x) implies (u-u2)x = uv

                          implies x= v/(1-u)       (u not equal to 0)

      therefore,for x to be real u not equals to 1.

hence v can take any real value,and u can take all real values except 1&0.

ans (b)